Determine the frequencies of the fundamental mode and the next two harmonics

that can be set up on a sitar string of length 1.0 m. Take the speed of waves of the

string to be 2.8 *103ms-1.


1
Expert's answer
2021-05-03T10:28:41-0400

"f_1=\frac {v}{2L}=1.4~kHz,"

"f_2=2\cdot \frac{v}{2L}=2f_1=2.8~kHz,"

"f_3=3\cdot \frac{v}{2L}=3f_1=4.2~kHz."